# Convergence of the extended Kalman filter with small and state-dependent noise

### v1

@@ 3.1 Extended Kalman filter
$\dot{Q}(t)$ $=\big[\nabla_{y}f-g\ell^{\top}(\ell\ell^{\top})^{-1}\nabla_{y}h\big](t,M(t))Q(t)+Q(t)\big[\nabla_{y}f-g\ell^{\top}(\ell\ell^{\top})^{-1}\nabla_{y}h\big]^{\top}(t,M(t))$
$~~~~~~-Q(t)[\nabla_{y}h^{\top}(\ell\ell^{\top})^{-1}\nabla_{y}h](t,M(t))Q(t)+a(t,M(t)),$ (31)
where
$a(t,M(t))=\sigma\sigma^{\top}(t,M(t))+g(I-\ell^{\top}(\ell\ell^{\top})^{-1}\ell)g^{\top}(t,M(t)).$ (32)
@@ 3.2 Some definitions
We will now introduce some essential concepts for this study. We will use the expression a "family of" function or process when this function or process depends on $\varepsilon$. Furthermore, we fix a family of filtered probability spaces, denoted by $\left(\Omega,\pazocal{F},\mathbb{F},\mathbb{P}\right)$ and we assume that the family of functions $f,~h,~\sigma,~g,$$\ell$ are measurable and locally bounded. The filtration, denoted by $\mathbb{F}$, is such that all the processes defined in the paper are adapted with respect to it.
Definition 3.2(Observable Process)
Let us denote by $\mathbb{F}^{Z}$ the observable filtration generated by the observation process $(Z(t))_{t\geq 0}$. We will call an observable process any $\mathbb{F}^{Z}$-predictable process.
This means that any observable process is adapted with respect to the natural filtration $\mathbb{F}^{Z}$.
Definition 3.3
@@
    $|\varphi(t,x_{1})-\varphi(t,x_{2})-F(t)(x_{1}-x_{2})|\leq\mu^{\varepsilon}|x_{1}-x_{2}|,$ (33)
    for some family of numbers $\mu^{\varepsilon}$ converging to zero as $\varepsilon\to 0$; the process $F$ will be called an almost derivative of $f$.
@@
Definition 3.5
A family of process $(\pazocal{X}^{\varepsilon}(t))_{t\geq 0}$ is said to be bounded in $L^{\infty-}$ if for any $q\in(0,\infty)$, there exists $\varepsilon_{q}>0$ such that $|\pazocal{X}^{\varepsilon}(t)|_{q}$ is bounded uniformly in $t\geq 0$ and $0<\varepsilon<\varepsilon_{q}$.
Definition 3.6(Stochastic Process of Order $\varepsilon^{\alpha}$)
A family of stochastic processes $(\pazocal{X}^{\varepsilon}(t))_{t\geq 0}$, $\varepsilon>0$ is said to be of order $\varepsilon^{\alpha}$ for some $\alpha>0$ if for any $q\in[1,\infty)$ there exist constants $\varepsilon_{q}>0,~C_{q}>0$, such that
$\varepsilon^{-\alpha}\left|\pazocal{X}^{\varepsilon}(t)\right|_{q}\leq C_{q}~~\text{for all}~~(t,\varepsilon)\in[0,\infty)\times(0,\varepsilon_{q}).$
Similarly, a family of random variables $\pazocal{X}^{\varepsilon}$, $\varepsilon>0$ is said to be of order $\varepsilon^{\alpha}$ for some $\alpha>0$ if for any $q\in[1,\infty)$ there exist constants $\varepsilon_{q}>0,~C_{q}>0$, such that
$\varepsilon^{-\alpha}\left|\pazocal{X}^{\varepsilon}\right|_{q}\leq C_{q}~~\text{for all}~~\varepsilon\in(0,\varepsilon_{q}).$
@@
$\dot{\zeta}(t)=A(t)\zeta(t),~~~\zeta(0)=I.$ (35)
We will say that $(A(t))_{t\geq 0}$ is exponentially stable if there exist some positive constant numbers $C$ and $c$ such that for $s\leq t$,
$|\zeta(t)|\leq Ce^{(-c(t-s))}|\zeta(s)|.$ (36)
Remark 3.8
In case of constant matrix $A$, we can show that $A$ is exponentially stable if and only if its eigenvalues have negative real part.
Definition 3.9
@@
Definition 3.10
Consider a family $(\pazocal{K}(t))_{t\geq 0}$ of absolutely continuous adapted (w.r.t. the observable filtration) processes with values in the class of symmetric positive definite $n\times n$ matrices and a family $(k(t))_{t\geq 0}$ of locally bounded deterministic functions with positive values. We will say that a family $(A(t))_{t\geq 0}$ of measurable locally bounded process with values in the class $n\times n$ matrices, is $(\pazocal{K}(t),k(t))_{t\geq 0}$-stable if for all $t\geq 0$,
$\dot{\pazocal{K}}(t)\succeq A(t)\pazocal{K}(t)+\pazocal{K}(t)A^{\top}(t)+k(t)\pazocal{K}(t).$ (37)
Remark 3.11
The link between these two stability concepts is given below by
Lemma 5.3.
@@ 4 Error estimation for the extended Kalman filter
When estimating a hidden signal, it is of relevance to know whether the scale of the error between the estimate and the true signal is large or not. This is why we are interested in conditions under which the extended Kalman filter for problem (1) provides a good approximation in the case of a strongly injective drift coefficient of the observation process. We show in the following theorem that the strongly injectivity combined with an initial guess of order $\sqrt{\varepsilon}$ and a set of technical conditions ensure that the estimation error Y(t)M(t)Y(t)-M(t) is of order $\sqrt{\varepsilon}$ at any future time point $t$.
Theorem 4.1
Let $(M(t),Q^{\varepsilon}(t))_{t\geq 0}$ be an extended Kalman filter for problem (1) such that $M$ and $Q=\frac{1}{\varepsilon}Q^{\varepsilon}$ satisfy (29) and (31) respectively, and in addition
@@
    $a=\sigma\sigma^{\top}-g(I-\ell^{\top}(\ell\ell^{\top})^{-1}\ell)g^{\top}$ and $(\ell\ell^{\top})^{-1}$ are uniformly elliptic,
    and the quotient between the largest and the smallest eigenvalues of $Q(0)$ is bounded.
@@
Remark 4.2
Assumptions (A2)-(A4) of this theorem can be checked for the $SI^{\pm}S$ epidemic model with constant population size. Moreover, all compartment sizes are assumed strictly positive. The assumption (A2) is fulfilled because of the strictly positive compartment sizes. The hidden signal and observation drift coefficients are respectively,
$f(Y(\tau))=$ $~~\beta(1-(X_{1}(0)+\varepsilon Y(\tau))-(X_{2}(0)+\varepsilon Z(\tau)))(X_{1}(0)+\varepsilon Y(\tau))$-
$~~(\alpha+\rho^{-})(X_{1}(0)+\varepsilon Y(\tau)),$
$h(Y(\tau))=$ $~~\alpha(X_{1}(0)+\varepsilon Y(\tau))-\rho^{+}(X_{1}(0)+\varepsilon Y(\tau)).$
These functions are $\pazocal{C}^{1}$ in $Y$ and $h$ is almost linear since it is linear, hence (A3) is fulfilled. The function $f$ is also almost linear with almost derivative the coefficient of the linear term in $f$, which is
$\beta\varepsilon\left(1-2X_{1}(0)-(X_{2}(0)+\varepsilon Z(\tau))\right)-(\alpha+\rho^{-})\varepsilon.$
For assumption (A4), $h$ is strongly injective since it is linear, and $a$ and $(\ell\ell^{\top})^{-1}$ are strictly positive scalars since all compartments sizes are strictly positive. The initial matrix $Q(0)$ is a scalar and can be chosen to satisfy the assumption.
For the first assumption, the scaled initial error estimate $Q^{-1/2}(0)(Y(0)-M(0))$ should be of order $\sqrt{\varepsilon}$. This implies that the accuracy of the scaled initial estimate of absolute population size should be of order $\sqrt[4]{N}$ since $\varepsilon=\dfrac{1}{\sqrt{N}}$.
@@
Lemma 4.3
Let $(M(t),Q^{\varepsilon}(t))_{t\geq 0}$ be an EKF satisfying assumptions (A1)-(A3) of Theorem 4.1. Suppose also that
$Q(t)\big(\nabla_{y}h^{\top}\nabla_{y}h\big)(t,M(t))Q(t)+a(t,M(t))\succeq k(t)Q(t),$ (38)
for some family of deterministic positive functions $k(t)$ such that
$Q(t)+Q^{-1}(t)\preceq Ck(t)I.$ (39)
Then $Q^{-\frac{1}{2}}(t)\big(Y(t)-M(t)\big)$ is of order $\sqrt{\varepsilon}$.
The proof of this Lemma is based on the following lemma, which is proven in picard1991efficiency.
Lemma 4.4
Let $\widetilde{W}(t)$ be a $\mathbb{F}$-Brownian motion with values in $\mathbb{R}^{r}$, $n\in\mathbb{N}$, and let $A(t)$ be a family of $\mathbb{F}$- adapted $(\pazocal{K}(t),k(t))$-stable processes where $\pazocal{K}(t)$ is $\mathbb{F}$ adapted. We suppose that $\pazocal{X}(t)$ is a family of $\mathbb{R}^{n}$-valued semi-martingales satisfying
$d\pazocal{X}(t)=A(t)\pazocal{X}(t)dt+F(t)dt+D(t)d\widetilde{W}(t),$ (40)
where $F(t)$ and $D(t)$ are predictable processes satisfying
$\pazocal{X}^{\top}(t)\pazocal{K}^{-1}(t)F(t)$ $\leq~\alpha k(t)\pazocal{X}^{\top}(t)\pazocal{K}^{-1}(t)\pazocal{X}(t)+\pazocal{O}(k(t)),$ (41)
$D^{\top}(t)\pazocal{K}^{-1}(t)D(t)$ $=~\pazocal{O}(k(t)),$ (42)
for some $\alpha\in(0,\frac{1}{2})$.
If $\pazocal{X}^{\top}(0)\pazocal{K}^{-1}(0)\pazocal{X}(0)$ is bounded in $L^{q}$, then for any $q\geq 1$ and $\varepsilon$ small enough, the process $\pazocal{X}^{\top}(t)\pazocal{K}^{-1}(t)\pazocal{X}(t)$ is bounded in $L^{q}$.
Proof
of Lemma 4.3
Let us consider $\pazocal{K}(t)=\varepsilon Q(t)$ and let us recall the dynamics of our signal $Y$ and of the extended Kalman filter $M$ from (1) and (29), respectively.
@@
$+\left[f(t,Y(t))-f(t,M(t))-\nabla_{y}f(t,M(t))\pazocal{X}(t)\right$.
$\left.-\Big(g\ell^{\top}(t,M(t))+Q(t)\nabla_{y}h^{\top}(t,M(t))\Big)(\ell\ell^{\top})^{-1}\right$.
$\times\left.\Big(h(t,Y(t))-(h(t,M(t))-\nabla_{y}h(t,M(t))\pazocal{X}(t))\Big)\right]dt+\sqrt{\varepsilon}\Big(\sigma(t,Y(t)),$
@@
We are then in the setting of Lemma 4.4 with
@@
$A(t)~:=$ $\nabla_{y}f(t,M(t))-\big(g\ell^{\top}(t,M(t))+Q(t)\nabla_{y}h^{\top}\big)(\ell\ell^{\top})^{-1}\nabla_{y}h(t,M(t)),$
$F(t)~:=$ $\left[f(t,Y(t))-f(t,M(t))-\nabla_{y}f(t,M(t))\pazocal{X}(t)-\Big(g\ell^{\top}(t,M(t))+Q(t)\nabla_{y}h^{\top}(t,M(t))\Big)\right$.
$\times(\ell\ell^{\top})^{-1}(t,M(t))\left.\Big(h(t,Y(t))-(h(t,M(t))-\nabla_{y}h(t,M(t))\pazocal{X}(t))\Big)\right],$
$D(t)~:=$ $\sqrt{\varepsilon}\Big(\sigma(t,Y(t)),~g(t,Y(t))-\big[g\ell^{\top}(t,M(t))+Q(t)\nabla_{y}h^{\top}(t,M(t))\big]\big(\ell\ell^{\top}\big)^{-1}\ell(t,Y(t))\Big),$
$\widetilde{W}(t):=$~ $\left(\begin{array}[]{c}W^{(1)}(t)\\ W^{(2)}(t)\end{array}\right).$
By definition, for any $k(t)$, saying that $A(t)$ is $(Q(t),k(t))$-stable means that
$\dot{Q}(t)$ $=\Big[\nabla_{y}f-g\ell^{\top}(\ell\ell^{\top})^{-1}\nabla_{y}h\Big]Q(t)+Q(t)\Big[\nabla_{y}f-g\ell^{\top}(\ell\ell^{\top})^{-1}\nabla_{y}h\Big]$
$~~~~~-Q(t)\nabla_{y}h^{\top}(\ell\ell^{\top})^{-1}\nabla_{y}hQ(t)+a(t,M(t))$
$\succeq\Big[\nabla_{y}f-g\ell^{\top}(\ell\ell^{\top})^{-1}\nabla_{y}h\Big]Q(t)-Q(t)\nabla_{y}h^{\top}(\ell\ell^{\top})^{-1}\nabla_{y}hQ(t)$
$~~~~~+Q(t)\Big[\nabla_{y}f-g\ell^{\top}(\ell\ell^{\top})^{-1}\nabla_{y}h\Big]^{\top}-Q(t)\nabla_{y}h^{\top}(\ell\ell^{\top})^{-1}\nabla_{y}hQ(t)+k(t)Q(t),$
and this implies that $a(t,M(t))\succeq-Q(t)\nabla_{y}h^{\top}(\ell\ell^{\top})^{-1}\nabla_{y}hQ(t)+k(t)Q(t)$, which is equivalent to the condition (38). Now, we have to check the two conditions (41) and (42).
Condition (42) can be deduced from the assumption (39) as follows. We have that
$Q(t)+Q^{-1}(t)\preceq Ck(t)I$ This implies that
$D^{\top}(t)(Q(t)+Q^{-1}(t))D(t)$ $\preceq D^{\top}(t)Ck(t)ID(t)$ (43)
$\text{and thus }~~D^{\top}(t)Q(t)D(t)+D^{\top}(t)Q^{-1}(t)D(t)$ $\preceq Ck(t)|D(t)|^{2}.$ (44)
Since $Q(t)$ and $Q^{-1}(t)$ are positive definite, hence $D^{\top}(t)Q(t)D(t)\succeq 0$ and
$D^{\top}(t)Q^{-1}(t)D(t)\succeq 0$. In addition, $|D(t)|$ is bounded, then $D^{\top}(t)Q^{-1}(t)D(t)=O(k(t))$.
Now, let us prove the first condition. Instead of the first condition, we are going to show that
$F^{\top}(t)\pazocal{K}^{-1}(t)F(t)\leq\alpha_{\epsilon}k^{2}(t)\pazocal{X}^{\top}(t)\pazocal{K}^{-1}(t)\pazocal{X}(t),$ (45)
which will imply that
$\pazocal{X}^{\top}(t)\pazocal{K}^{-1}(t)F(t)\leq\alpha^{1/2}_{\epsilon}k(t)\pazocal{X}^{\top}(t)\pazocal{K}^{-1}(t)\pazocal{X}(t),$ (46)
according to the Remark 4.5 below.
We have
@@
Thus, since $f$ and $h$ are almost linear, we obtain the following estimate,
$F^{\top}(t)Q^{-1}(t)F(t)=$ $~B^{\top}Q^{-1}B-H^{\top}(\ell\ell^{\top})^{-1}(g(t,M(t))\ell^{\top}$
$+Q(t)\nabla_{y}h(t,M(t)))^{\top}(\ell\ell^{\top})^{-1}Q^{-1}B$

### v2

@@ 3.1 Extended Kalman filter
$\begin{split}\dot{Q}(t)=&-Q(t)[\nabla_{y}h^{\top}(\ell\ell^{\top})^{-1}\nabla_{y}h](t,M(t))Q(t)\\ &+\big[\nabla_{y}f-g\ell^{\top}(\ell\ell^{\top})^{-1}\nabla_{y}h\big](t,M(t))Q(t)\\ &+Q(t)\big[\nabla_{y}f-g\ell^{\top}(\ell\ell^{\top})^{-1}\nabla_{y}h\big]^{\top}(t,M(t))+{\Phi}(t,M(t)),\end{split}$ (3.11)
where the absolute term $\Phi$} is given by
${\Phi}(t,M(t))=\sigma\sigma^{\top}(t,M(t))+g(I-\ell^{\top}(\ell\ell^{\top})^{-1}\ell)g^{\top}(t,M(t)).\hskip 50.00008pt$ (3.12)
@@ 3.2 Some definitions
We will now introduce some key concepts from this study and recall the notation of the underlying filtered probability space $\left(\Omega,\pazocal{F},\mathbb{F},\mathbb{P}\right)$. The filtration $\mathbb{F}$ is such that all processes defined in this article are adapted with respect to it. We will use the expression a “family of” function or process when this function or process depends on $\varepsilon$. It is assumed that the family of functions $f,~h,~\sigma,~g,$$\ell$ are measurable and locally bounded. Throughout the paper, for a vector $x$, we denote by $|x|$ the Euclidean norm of $x$, and for a matrix $A$, the matrix norm $|A|$ is considered to be the supremum of $|Ax|$ over unit vectors $x$.
Definition 3.3 (Observable Process).
Let $\mathbb{F}^{Z}\subset\mathbb{F}$ be the observable filtration generated by the observation process $(Z(t))_{t\geq 0}$. We will call an observable process any $\mathbb{F}^{Z}$-adapted process.
From the above definition, it follows that any observable process is adapted with respect to the natural filtration $\mathbb{F}^{Z}$.
Definition 3.4.
@@
    $|\varphi(t,x_{1})-\varphi(t,x_{2})-F(t)(x_{1}-x_{2})|\leq\mu^{\varepsilon}|x_{1}-x_{2}|,$ (3.13)
    for some family of numbers $\mu=(\mu^{\varepsilon})_{\varepsilon>0}$ converging to zero as $\varepsilon\to 0$.
    The process $F$ will be called an almost derivative of $f$.
@@
Definition 3.6.
A family of processes $(\pazocal{X}^{\varepsilon}(t))_{t\geq 0}$ is said to be bounded in $L^{\infty-}$ if for any $q\in(0,\infty)$, there exists $\varepsilon_{q}>0$ such that $|\pazocal{X}^{\varepsilon}(t)|_{q}$ is bounded uniformly in $t\geq 0$ and $0<\varepsilon<\varepsilon_{q}$.
Definition 3.7 (Stochastic Process of Order $\varepsilon^{\kappa}$).
A family of stochastic processes $\pazocal{X}(t)=(\pazocal{X}^{\varepsilon}(t))_{t\geq 0}$, $\varepsilon>0$ is said to be of order $\varepsilon^{\kappa}$ for some $\kappa>0$ if for any $q\in[1,\infty)$ there exist constants $\varepsilon_{q}>0,~C_{q}>0$, such that
$\varepsilon^{-\kappa}\left|\pazocal{X}^{\varepsilon}(t)\right|_{q}\leq C_{q}~~\text{for all}~~(t,\varepsilon)\in[0,\infty)\times(0,\varepsilon_{q}).$
Similarly, a family of random variables $X=(X^{\varepsilon})$, $\varepsilon>0$ is said to be of order $\varepsilon^{\kappa}$ for some $\kappa>0$ if for any $q\in[1,\infty)$ there exist constants $\varepsilon_{q}>0,~C_{q}>0$, such that
$\varepsilon^{-\kappa}\left|{X^{\varepsilon}}\right|_{q}\leq C_{q}~~\text{for all}~~\varepsilon\in(0,\varepsilon_{q}).$
Remark 3.8.
To avoid cluttering the notation, we usually suppress the superscript index $\varepsilon$ in the notation of families of function, stochastic processes, and random variables, but keep this dependency in mind.
@@
$\dot{\zeta}(t)=A(t)\zeta(t),~~~\zeta(0)=I.$ (3.15)
We will say that $(A(t))_{t\geq 0}$ is exponentially stable if there exist some positive constants $C$ and $c$ such that for $s\leq t$,
$|\zeta(t)|\leq Ce^{(-c(t-s))}|\zeta(s)|.$ (3.16)
Remark 3.10.
In case of constant matrix $A$, it is well-known that $A$ is exponentially stable if and only if its eigenvalues have negative real part.
Definition 3.11.
@@
Definition 3.12.
Consider a family $(\pazocal{K}(t))_{t\geq 0}$ of absolutely continuous adapted (w.r.t. the observable filtration) processes with values in the class of symmetric positive definite $n\times n$ matrices and a family $(k(t))_{t\geq 0}$ of locally bounded deterministic functions with positive values. We will say that a family $(A(t))_{t\geq 0}$ of measurable locally bounded process with values in the class $n\times n$ matrices, is $(\pazocal{K}(t),k(t))_{t\geq 0}$-stable, or in short, that $A$ is $(\pazocal{K},k)$-stable, if for all $t\geq 0$,
$\dot{\pazocal{K}}(t)\succeq A(t)\pazocal{K}(t)+\pazocal{K}(t)A^{\top}(t)+k(t)\pazocal{K}(t).$ (3.17)
Note that the relationship between these two stability concepts will be established below in Lemma 5.2.
@@ 4 Error estimation for the extended Kalman filter
When estimating a hidden signal, it is of relevance to know whether the scale of the error between the estimate and the true signal is large or not. This is why we are interested in conditions under which the extended Kalman filter for problem (1.1) provides a good approximation in the case of a strongly injective drift coefficient of the observation process. We show in the following theorem that the strongly injectivity combined with an initial guess of order $\sqrt{\varepsilon}$ and a set of technical conditions ensure that the estimation error Y(t)M(t)Y(t)-M(t) is of order $\sqrt{\varepsilon}$ at any future time point $t$.
Theorem 4.1.
Let $(M(t),Q^{\varepsilon}(t))_{t\geq 0}$ be an extended Kalman filter for problem (1.1) such that $M$ and $Q=\frac{1}{\varepsilon}Q^{\varepsilon}$ satisfy (3.9) and (3.11) respectively, and in addition
@@
    ${\Phi}=\sigma\sigma^{\top}-g(I-\ell^{\top}(\ell\ell^{\top})^{-1}\ell)g^{\top}$ and $(\ell\ell^{\top})^{-1}$ are uniformly elliptic, and
    the quotient between the largest and the smallest eigenvalues of $Q(0)$ is bounded.
@@
Remark 4.2.
Assumptions (A2)-(A4) of this theorem can be checked for the $SI^{\pm}S$ epidemic model considered in Sec. 2 on each small interval $t_{i},t_{i+1}),~i=0,\ldots,N_{t}-1$, under the natural assumption, that all compartment sizes are strictly positive, i.e., the proportions $Y$ and $Z$ of nondetected and detected infected take values only in the open interval $(0,1)$. The assumption (A2)) is fulfilled by the definition of the functions $\sigma$, $g$ and $\ell$ which are bounded for $y,z\in(0,1)$. The drift coefficients $f$ and $h$ in the SDEs for the hidden signal and the observation are linear functions of $y$, since the proportion of susceptible individuals $s$ is treated as a constant. Thus $f$ and $h$ are $\pazocal{C}^{1}$ functions, and also almost linear, hence (A3) is fulfilled.
For assumption (A4) it can be observed, that $h$ is strongly injective since it is linear, and $\Phi$} and $(\ell\ell^{\top})^{-1}$ are strictly positive scalars since all compartment sizes are strictly positive. The initial matrix $Q(0)$ is a scalar and can be chosen to satisfy the assumption.
@@
Lemma 4.3.
Let $(M(t),Q^{\varepsilon}(t))_{t\geq 0}$ be an EKF satisfying assumptions (A1)-(A4) of Theorem 4.1. Suppose also that $Q=\varepsilon^{-1}Q^{\varepsilon}$ satisfies
$Q(t)\big(\nabla_{y}h^{\top}(\ell\ell^{\top})^{-1}\nabla_{y}h\big)(t,M(t))Q(t)+{\Phi}(t,M(t))\succeq k(t)Q(t),$ (4.1)
for some family of deterministic positive functions $k(t)$ and the function $\Phi$} given in (A4). Further, it is assumed that there is a constant $C>0$ such that
$Q(t)+Q^{-1}(t)\preceq Ck(t)I.$ (4.2)
Then $Q^{-1/2}(t)\big(Y(t)-M(t)\big)$ is of order $\sqrt{\varepsilon}$.
The proof of this Lemma is based on the following lemma, which is proven in [22].
Lemma 4.4.
Let $\widetilde{W}(t)$ be a $\mathbb{F}$-Brownian motion with values in $\mathbb{R}^{r}$, $n\in\mathbb{N}$, and let $A$ be a family of $\mathbb{F}$-adapted $(\pazocal{K},k)$-stable processes where $\pazocal{K}$ is $\mathbb{F}$-adapted and $k$ is some family of deterministic positive and locally bounded functions. We suppose that $\pazocal{X}(t)$ is a family of $\mathbb{R}^{n}$-valued semi-martingales satisfying
$d\pazocal{X}(t)=A(t)\pazocal{X}(t)dt+F(t)dt+D(t)d\widetilde{W}(t),$ (4.3)
where $F(t)$ and $D(t)$ are $\mathbb{F}$-adapted processes satisfying for some $\eta\in[0,1/2)$
$\pazocal{X}^{\top}(t)\pazocal{K}^{-1}(t)F(t)$ $\leq~\eta k(t)\pazocal{X}^{\top}(t)\pazocal{K}^{-1}(t)\pazocal{X}(t)+\pazocal{O}(k(t)),$ (4.4)
$D^{\top}(t)\pazocal{K}^{-1}(t)D(t)$ $=~\pazocal{O}(k(t)).$ (4.5)
If $\pazocal{X}^{\top}(0)\pazocal{K}^{-1}(0)\pazocal{X}(0)$ is bounded in $L^{q}$, then for any $q\geq 1$ and $\varepsilon$ small enough, the process
$\pazocal{X}^{\top}(t)\pazocal{K}^{-1}(t)\pazocal{X}(t)$ is bounded in $L^{q}$.
Proof.
(of Lemma 4.3) Let us recall the dynamics of the signal $Y$ and of the extended Kalman filter $M$ from (1.1) and (3.9), respectively.
@@
$+\Big[f(t,Y(t))-f(t,M(t))-\nabla_{y}f(t,M(t))\pazocal{X}(t)$
$-\big(g\ell^{\top}(t,M(t))+Q(t)\nabla_{y}h^{\top}(t,M(t))\big)(\ell\ell^{\top})^{-1}$
$\times\big(h(t,Y(t))-(h(t,M(t))-\nabla_{y}h(t,M(t))\pazocal{X}(t))\big)\Big]dt+\sqrt{\varepsilon}\Big(\sigma(t,Y(t)),$
@@
Application of Lemma 4.4
The above SDE for $\pazocal{X}$ can be rewritten in terms of the notation used in Lemma 4.4 by using following settings
$\pazocal{K}(t)$~ $=Q(t),$
@@
$A(t)~:=$ $\nabla_{y}f(t,M(t))-\big(g\ell^{\top}(t,M(t))+Q(t)\nabla_{y}h^{\top}\big)(\ell\ell^{\top})^{-1}\nabla_{y}h(t,M(t)),$ (4.6)
$F(t)~:=$ $f(t,Y(t))-f(t,M(t))-\nabla_{y}f(t,M(t))\pazocal{X}(t)$ (4.7)
$-\big(g\ell^{\top}(t,M(t))+Q(t)\nabla_{y}h^{\top}(t,M(t))\big)$ (4.8)
$\times(\ell\ell^{\top})^{-1}(t,M(t))\big(h(t,Y(t))-(h(t,M(t))-\nabla_{y}h(t,M(t))\pazocal{X}(t))\big)$ (4.9)
$D(t)~:=$ $\sqrt{\varepsilon}\big(\sigma(t,Y(t)),~g(t,Y(t))$ (4.10)
$-\big[g\ell^{\top}(t,M(t))+Q(t)\nabla_{y}h^{\top}(t,M(t))\big]\big(\ell\ell^{\top}\big)^{-1}\ell(t,Y(t))\big),$ (4.11)
$\widetilde{W}(t):=$~ $\left(\begin{array}[]{c}W^{(1)}(t)\\ W^{(2)}(t)\end{array}\right).$ (4.14)
We will now verify the assumptions from Lemma 4.4.
AA is $(Q,k)$-stable
We recall the Riccati ODE (3.11) for $Q$
$\dot{Q}(t)=$ $-Q(t)\nabla_{y}h^{\top}(\ell\ell^{\top})^{-1}\nabla_{y}hQ(t)+\big[\nabla_{y}f-g\ell^{\top}(\ell\ell^{\top})^{-1}\nabla_{y}h\big]Q(t)$
$+Q(t)\big[\nabla_{y}f-g\ell^{\top}(\ell\ell^{\top})^{-1}\nabla_{y}h\big]^{\top}+{\Phi}(t,M(t)),$
and condition (4.1) which implies that for some family of deterministic positive functions $k$ it holds ${\Phi}(t,M(t))\succeq-Q(t)\nabla_{y}h^{\top}{(\ell\ell^{\top})^{-1}}\nabla_{y}hQ(t)+k(t)Q(t)$. Substituting this in the above ODE and using the definition of $A$, see above in (4.6), yields
$\dot{Q}(t)\succeq$ $-2Q(t)\nabla_{y}h^{\top}(\ell\ell^{\top})^{-1}\nabla_{y}hQ(t)+\Big[\nabla_{y}f-g\ell^{\top}(\ell\ell^{\top})^{-1}\nabla_{y}h\Big]Q(t)$
$+Q(t)\Big[\nabla_{y}f-g\ell^{\top}(\ell\ell^{\top})^{-1}\nabla_{y}h\Big]^{\top}+k(t)Q(t),$
$=$ $A(t)Q(t)+Q(t)A^{\top}(t)+k(t)Q(t).\qquad$
This proves that $A$ is $(Q,k)$-stable.
Verfication of condition (4.4)
For that, it is sufficient to show that for some family of functions $\overline{\eta}$ with values in [0,1/4)[0,{1}/{4}), it holds
$F^{\top}(t)\pazocal{K}^{-1}(t)F(t)\leq\overline{\eta}k^{2}(t)\pazocal{X}^{\top}(t)\pazocal{K}^{-1}(t)\pazocal{X}(t).$ (4.15)
This follows from an application of the Cauchy-Schwarz inequality corresponding to the inner product defined by the positive definite matrix $\pazocal{K}^{-1}(t)$ to
$\left|\pazocal{X}^{\top}(t)\pazocal{K}^{-1}(t)F(t)\right|$ $\leq\Big(\pazocal{X}^{\top}(t)\pazocal{K}^{-1}(t)\pazocal{X}(t)\Big)^{1/2}\Big(F^{\top}(t)\pazocal{K}^{-1}(t)F(t)\Big)^{1/2}$
$\leq\Big(\pazocal{X}^{\top}(t)\pazocal{K}^{-1}(t)\pazocal{X}(t)\Big)^{1/2}\Big(\overline{\eta}k^{2}(t)\pazocal{X}^{\top}(t)\pazocal{K}^{-1}(t)\pazocal{X}(t)\Big)^{1/2}$
$=(\overline{\eta})^{1/2}k(t)\pazocal{X}^{\top}(t)\pazocal{K}^{-1}(t)\pazocal{X}(t).$
Hence, we have the relation $\eta=\overline{\eta})^{1/2}$. Next, we have for $\pazocal{K}(t)=Q(t)$
@@
Thus, since $f$ and $h$ are almost linear, there exists a family of positive constants $\overline{\mu}=(\overline{\mu}^{\varepsilon})_{\varepsilon>0}$ converging to 0 as $\varepsilon\rightarrow 0$ , such that $|B|,|H|\leq\overline{\mu}|Y(t)-M(t)|$. Moreover, we have that,
$F^{\top}(t)Q^{-1}(t)F(t)=$ $~B^{\top}Q^{-1}B$
$-H^{\top}(\ell\ell^{\top})^{-1}\Big(g(t,M(t))\ell^{\top}+Q(t)\nabla_{y}h(t,M(t))\Big)^{\top}(\ell\ell^{\top})^{-1}Q^{-1}B$